Opening the book…
After completing this section, you should be able to:
- Determine odds from probabilities.
- Determine probabilities from odds.
- Calculate the expected value of an experiment.
- Interpret the expected value of an experiment.
- Use expected value to analyze applications.
A particular lottery instant-win game has \(2\) million tickets available. Of those, \(500,000\) win a prize. If there are \(500,000\) winners, then it follows that there are \(1,500,000\) losing tickets. When we evaluate the risk associated with a game like this, it can be useful to compare the number of ways to win the game to the number of ways to lose. In the case of this game, we would compare the \(500,000\) wins to the \(1,500,000\) losses. In other words, there are three losing tickets for every winning ticket. Comparisons of this type are the focus of this section.
Computing Odds
The ratio of the number of equally likely outcomes in an event, \(E\), to the number of equally likely outcomes not in the event, \(E^{'}\), odds for (or odds in favor of) the event. The opposite ratio (the number of outcomes not in the event to the number in the event to the number in the event is called the odds against the event. The formula is given below
Let’s say, based on past data, in a certain city, it rains \(60\) days in a year, and it doesn't rain \(305\) days a year. Find the odds in favor of a rainy day and interpret your answer.
Now, suppose you're trying to find the odds in favor of a rainy day:
Favorable outcomes (rainy days) = \( 60\)
Unfavorable outcomes (non-rainy days) = \(305\)
\[\text{Odds in favor of rain} = 60{:}305=\frac{60}{305}=\frac{12}{61}=12{:}61\nonumber\]
Interpretation: The odds in favor of rain on a randomly chosen day are \(12{:}61,\) meaning for every \( 12\) rainy day, there are \( 61\) days without rain.
For an event \(E\)
\begin{align*}
\text{P(not E )} &= 1 - P(E) \\
\text{Odd in favor of E } &= \dfrac{P(E)}{P(\text{not}~E)}\\
\text{Odd against E } &= \dfrac{P(\text{not} ~E)}{P(E)}
\end{align*}
Figure \(\PageIndex{2}\): Understanding Odd in Favor and Odd AgainstIf we know the numbers of favorable and unfavorable outcomes, we can use the following to calculate the odds against and in favor.
| Type | Formula | Interpretation |
|---|---|---|
| Odds in Favor | \[\text{~Odd in Favor}=\dfrac{~\text{Number of favorable outcomes}}{~\text{Number of unfavorable outcomes}}\nonumber\] |
It compares how often an event happens vs. how often it does not. |
| Odds Against | \[\text{~Odd Against}=\dfrac{~\text{Number of unfavorable outcomes}}{~\text{Number of favorable outcomes}}\nonumber\] | It compares how often an event does not happen vs. how often it does. |
Both odds and probabilities are calculated as ratios. To avoid confusion, we will always use fractions, decimals, or percents for probabilities and colons to indicate odds. The rules for simplifying fractions also apply to odds. Thus, the odds for winning a prize in the game described in the section opener are \( 500,000{:}1,500,000=1{:}3\) and the odds against winning this prize are \(3{:}1\). These would often be expressed in words as “the odds of winning are one to three in favor” or “the odds of winning are three to one against.
Example \(\PageIndex{1}\): Computing Odds
If you draw a card at random from a standard deck, what are the odds of drawing a face card?
- Answer
-
One Method:
There are \(12\) of those outcomes in the event, E, “drawing a face card,” so E ={J♥, Q♥, K♥,J♥, Q♦, K♦, J♣, Q♣, K♣ J♠, Q♠, K♠}. Note that a deck has \(52\) cards.
Number of favorable outcomes (number of face cards) = \(12\)
Number of unfavorable outcomes (number of not face cards) = \(52-12=40\)
\[\text{~Odd in Favor}=\dfrac{~\text{Number of favorable outcomes}}{~\text{Number of unfavorable outcomes}}=\frac{12}{40}=\frac{3}{10}\nonumber\]
\[\text{~Odd against}=\dfrac{~\text{Number of unfavorable outcomes}}{~\text{Number of favorable outcomes}}=\frac{40}{12}=\frac{10}{3}\nonumber\]
Another Method:
We can find odds another way as shown in the following
\begin{aligned}
\text{P(E )} &= \frac{12}{52}\\&=\frac{3}{13}\\ \text{P(not E )} &= 1- \frac{3}{13}=\frac{10}{13}
\end{aligned}
\begin{aligned}
\text{Odd in favor of E } = \dfrac{P(E)}{P(\text{not} E)}= \dfrac{\frac{3}{13}}{\frac{10}{13}}=\frac{3}{13}\times \frac{13}{10} =\frac{3}{10} = 3{:}10
\end{aligned}
Example \(\PageIndex{2}\): Converting Probabilities to Odds
Given the following probabilities of an event, find the corresponding odds for and against that event.
\(\text{P(E )} = 17\%\)
- Answer
-
One Method:
Note that \(\text{P(E )} = 17\%\) means, out of \(100\), \(17\) times that outcome occurs. In other words,
Number of favorable outcomes for event E = \(17\)
Number of unfavorable outcomes for event E = \(100-17=83\)
\[\text{~Odd in Favor}=\dfrac{~\text{Number of favorable outcomes}}{~\text{Number of unfavorable outcomes}}=\frac{17}{83}\nonumber\]
\[\text{~Odd against}=\dfrac{~\text{Number of unfavorable outcomes}}{~\text{Number of favorable outcomes}}=\frac{83}{17}\nonumber\]
Another Method
We’ll use another formula:
\[\text{P(E )} = 0.17=\frac{17}{100}\nonumber\]
\[\text{P(not E )} = 1- 0.17 =0.83=\frac{83}{100}\nonumber\]
\begin{align*}
\text{Odd in favor of E } &= \dfrac{\frac{17}{100}}{\frac{83}{100}}\\&=\frac{17}{100}\times \frac{100}{83}\\&= \frac{17}{83}\\&=17{:}83
\end{align*}
Odds as a Ratio of Probabilities
We can use these formulas to convert between probabilities and odds.
We can also think of odds as a ratio of probabilities. Consider the instant-win game from the section opener again, with \(500,000\) winning tickets out of \(2,000,000\) total tickets.
Number of favorable tickets is \(500,000\).
Number of unfavorable tickets is \(2,000,000-500,000=1,500,000\).
If a player buys one ticket, the probability of winning is
\[ \frac{5000,000}{2,000,000}=\frac{1}{4}\nonumber\]
And the probability of losing is \(1- \frac{1}{4}=\frac{3}{4}\). Notice that the ratio of the probability of winning to the probability of losing is \(\frac{1}{4}\div\frac{3}{4}=\frac{1}{3}=1:3\), which matches the odds in favor of winning.
Here, "favorable" means the number of favorable outcomes.
If the odds in favor of \( E\) are \(\text{~favorable}{:}\text{~unfavorable},\) then the probability of \(E\) is
\[\text{~Total} =\text{~favorable}+\text{~unfavorable}\nonumber\]
\[\text{~P(E)} = \frac{~\text{favorable}}{\text{Total}}\nonumber\]
If the odds against of \( E\) are \(\text{~uavorable}{:}\text{~favorable},\) then the probability of \(E\) is
\[\text{~P(E)} = \frac{~\text{unfvorable}}{\text{Total}}\nonumber\]
Example \(\PageIndex{3}\): Converting Odds to Probability
Find \(P(E)\) if the odds in favor of \(E\) are \(4{:}9.\)
- Answer
-
The odds in favor of \(E\) are \(4 : 9\). That means, the number of favorable outcomes is \(4\), and the number of unfavorable outcomes is \(9\), and the number ot total outcomes is \(4+9=13\).
\[\text{~P(E)} = \frac{~\text{favorable}}{\text{~Total}}=\frac{4}{13}\nonumber\]
Some places, particularly state lottery websites, will use the words “odds” and “probability” interchangeably. Never assume that the word “odds” is being used correctly! Compute one of the odds/probabilities yourself to make sure you know how the word. Expected Valued is being used!
Expected Value
The casino game roulette has dozens of different bets that can be made. These bets have different probabilities of winning but also have different payouts. In general, the lower the probability of winning a bet, the more money a player wins for that bet. With so many options, is there one bet that’s “smarter” than the rest? What’s the best play to make at a roulette table? This section will develop the tools we need to answer these questions.
Many experiments have numbers associated with their outcomes. Some are easy to define; if you roll \(2\) dice, the sum of the numbers showing is a good example. In some card games, cards have different point values associated with them; for example, in some forms of the game rummy, aces are worth \(15\) points; \(10\) s, jacks, queens, and kings are worth \(10\); and all other cards are worth \(6\). The outcomes of casino and lottery games are all associated with an amount of money won or lost. These outcome values are used to find the expected value of an experiment: the mean of the values associated with the outcomes that we would observe over a large number of repetitions of the experiment.
That definition is a little vague; How many is “a large number?” In practice, it depends on the experiment; the number must be large enough that every outcome is expected to appear at least a few times. For example, if we’re rolling a standard 6-sided die and noting the number that shows, a few dozen replications should be sufficient to ensure the mean is representative. Since the probability of each outcome is \(\frac{1}{6}\), we would expect to see each outcome about \(8\) times over the course of \(48\) replications. However, if we’re talking about the Powerball lottery, where the probability of winning the jackpot is about \(\frac{1}{292,000,000}\), we would need several billion replications to ensure that every outcome appears a few times. Luckily, we can find the theoretical expected value before we even run the experiment for the first time.
If \(X\) represents the value of that outcome, \(P(X)\) represents the probability of the outcome. The expected value of the experiment is:
\[\Sigma (X)\times P(X))\nonumber\]
where \(\Sigma\) is the “sum,” meaning we add up the results of the formula that follows over all possible outcomes.
We can find the probability distribution table for the above diagram.
| Outcome and Value of the Outcome \((X)\) | Probability \(P(X)\) |
| Land on Red: \(\$5\) | \(\frac{1}{4}\) |
| Land on Blue: \(\$5\) | \(\frac{1}{4}\) |
| Land on Green: \(-\$5\) | \(\frac{1}{4}\) |
| Land on White:\(-\$1\) | \(\frac{1}{4}\) |
Key Points to Remember for Finding Expected Value
-
Identify all possible outcomes and their values.
These values can be positive or negative depending on the situation. -
Determine the probability of each outcome.
-
Set up a probability distribution table:
-
First column (or row) → outcome probabilities
-
Second column (or row) → corresponding outcome numerical value
-
-
Multiply each outcome value by its probability, then add all the products together to find the expected value.
Here are the most common areas where people use expected value:
Gambling and Games: Used to determine the average gain or loss in activities like lottery, cards, dice, and slot machines, and to decide if a game is fair or profitable.
Insurance: Insurance companies use expected value to set premiums by estimating the expected cost of accidents, health claims, or property damage. Finance and Investments: Used to evaluate risky investments and predict the average profit or loss over time.
Sports: Helps predict average points, wins, or scoring outcomes in a game.
Weather and Natural Events: Used to estimate expected rainfall, average temperature, or the likelihood of natural disasters.
Example \(\PageIndex{4}\): Finding Expected Values
Find the expected values of the following experiments.
An unfair six-sided die has sides numbered \(1,2,3,4,5,6\) with probabilities given in the table below. Find the expected value of one roll of the die.
| Value of Outcome | \(1\) | \(2\) | \(3\) | \(4\) | \(5\) | \(6\) |
|---|---|---|---|---|---|---|
| Probability | \(\frac{1}{8}\) | \(\frac{1}{8}\) | \(\frac{1}{8}\) | \(\frac{1}{4}\) | \(\frac{1}{4}\) | \(\frac{1}{8}\) |
- Answer
-
We multiply each value by its probability and sum the results.
\begin{align*}
\text{Expected value} &= 1\times\frac{1}{8}+2\times\frac{1}{8}+3\times\frac{1}{8}+4\times\frac{1}{4}+5\times\frac{1}{4}+ 6\times\frac{1}{8}\\
& =\frac{1}{8}+\frac{2}{8}+\frac{3}{8}+\frac{4}{4}+\frac{5}{4}+\frac{6}{8}\\
& =\frac{30}{8}=3.75
\end{align*}So, the expected value of one roll of the dice is \(3.750.\)
Let's look at another example. Where the probability distribution table must be created.
Example \(\PageIndex{5}\): Finding Expected Values
Find the expected values of the following experiments.
Draw a card from a well-shuffled standard deck of cards and note its rummy value (\(15\) for aces; \(10\) for tens, jacks, queens, and kings; \(5\) for everything else).
- Answer
-
Let’s find the value of the outcome and its corresponding probabilities. We have three outcomes that we are looking for:
- aces
- tens, jacks, queens, and kings
- Everything else.
Possible outcomes and probabilities for the given experiment Outcome Probability Value of Outcome \(\{A\}\) \(\frac{4}{52}\) \(15\) \(\{10, J, K, Q\}\) \(\frac{16}{52}\) \(10\) \(\{2, 3, 4, 5, 6, 7, 8, 9\}\) \(\frac{32}{52}\) \(5\) We multiply each value by its probability and add them.
\begin{align*}
\text{Expected value} &= 15\times\frac{4}{52}+10\times\frac{16}{52}+5\times\frac{32}{52}\\
& =\frac{60}{52}+\frac{160}{52}+\frac{160}{52}\\
& =\frac{380}{52}=7.3
\end{align*}
As we noted, the expected value of an experiment is the mean of the values we would observe if we repeated the experiment a large number of times. (This interpretation is due to an important theorem in the theory of probability called the Law of Large Numbers.)
Example \(\PageIndex{6}\): Finding Expected Value (Sum of Two Dice)
A bag contains \(3\) gold marbles, \(9\) silver marbles, and \(26\) black marbles. Someone offers to play this game: You randomly select one marble from the bag. If it is gold, you win \($4.00.\) If it is silver, you win \($3.00.\) If it is black, you lose \($1.00.\) What is your expected value if you play this game?
- Answer
To find the expected value, you need to determine the outcomes and their probabilities. Those are listed in the following table.
| Outcome | Value | Probability | Value \(\times\) probability |
|---|---|---|---|
| Gold | \($4\) | \(\frac{3}{38}\) | \(\frac{12}{38}\) |
| Silver | \($3\) | \(\frac{9}{38}\) | \(\frac{27}{38}\) |
| Black | \(-$1\) | \(\frac{26}{38}\) | \(\frac{-26}{38}\) |
\begin{aligned}
\text{Expected value} &= $\frac{12}{38}+$\frac{27}{38}+$\frac{-26}{38}\\
&= \frac{-4}{38}\\ &= - $0.3421
\end{aligned}
Example \(\PageIndex{7}\): Risk of Dying
A \(40\)-year-old man in the U.S. has a \(0.24\%\) risk of dying during the next year. An insurance company charges \(\$300\) annually for a life insurance policy that pays a \(\$100,000\) death benefit. What is the expected value for the person buying the insurance? Round your answer to the nearest dollar.
- Answer
-
To find the expected value, we need to determine the value of each outcome and its corresponding probability. Those are listed in the following table.
Outcomes and their Probabilities Outcomes (Events) Values \(n(O)\) Probability \(P(O)\) Survive \(- $300\) \(1- 0.0024 = 0.9976\) Die \($100,000- $300 = $99,700\) \(0.0024\) \[\text{Expected value} = (- $300)(0.9976) + ($99,700)(0.0024) =-$60\nonumber\]
Note: The negative sign in front of \($300\) is because we are looking for the insured person's expected value; if that person survives, she/he will lose \($300.\) You also see that we subtract \($300\) to find the value of the outcome die. It is because the insured person gets \($300\) less than the policy value because the insured person has already paid \($300\) for the policy.
Example \(\PageIndex{8}\): Every Game Has a Negative Expected Value to the player
Suppose someone offers to play a game with you. If you roll a die and get a \(6,\) you get \($10.\) However, getting a five or below means you lose \(\$1.\) Is this a game you’d want to play?
- Answer
-
Let’s look at the expected value: The probability of winning is \(\frac{1}{6}\), and the probability of losing is\(\frac{5}{6}\), so the expected value is
\[$10\times \frac{1}{6}+(-$1)\frac{5}{6}=-$0.83\nonumber\]That means, on average, you’ll come out ahead by about \(83\) cents every time you play this game.
It’s a great deal! On the other hand, if the winnings for rolling a \(6\) drop to \($3\), the expected value becomes
\[$3\times \frac{1}{6}+(-$1)\frac{5}{6}=-$0.33\nonumber\]This means you should expect to lose about \(33\) cents on average for every time you play.
Playing that game is not a good idea! This is how casinos and lottery corporations make money:
In the \(21\)st century, data analytics tools have revolutionized the way sports are coached and played. One tool in particular is used in football at crucial moments in the game. When a team faces a fourth down (the last possession in a series of four possessions, a fairly common occurrence), the coach faces a decision: Run one play to try to gain a certain number of yards, or kick the ball away to the other team. Here’s the interesting part of the decision: If the team “goes for it” and runs the play and they are successful, then they keep possession of the ball and can continue in their quest to score more points. If they are unsuccessful, then they lose possession of the ball, giving the other team an opportunity to score points. If, instead, the team punts, or kicks the ball away, then the other team gets possession of the ball, but in a worse position for them than if the original team goes for it and fails. To analyze this situation, data analysts have generated empirical probabilities for every fourth-down situation and computed the expected value (in terms of points) for each decision. Coaches frequently use those calculations when they decide which option to take!
In \(1654\), a French writer and amateur mathematician named Antoine Gombaud (who called himself the Chevalier du Mére) reached out to his gambling buddy Blaise Pascal to answer a question that he’d read about called the “problem of points.” The question goes like this: Suppose you’re playing a game that is scored using points, and the first person to earn \(5\) points is the winner. The game is interrupted with the score \(4\) points to \(2\). If the winner stood to win \($100\), how should the prize money be divided between the players? Certainly, the person who is \(1\) point away from victory should get more, but how much more?
We have developed tools in this section to answer this question. At its heart, it’s a question about conditional probabilities and expected value. Pascal first started thinking about it at the time, though those ideas hadn’t yet been invented. Pascal reached out to a colleague named Pierre de Fermat, and over the course of a couple of months, their correspondence with each other would eventually solve the problem. In the process, they first described conditional probabilities and expected values!
Pascal is remembered for the “arithmetical triangle” that is named after him (though he wasn’t the first person to discover it; see the section on the binomial distribution for more), as well as his work in geometry. In physics, Pascal worked on hydrodynamics and air pressure (the SI unit for pressure is named after him), and in philosophy, Pascal advocated a mathematical approach to philosophical problems.
Example \(\PageIndex{9}\): Using Expected Values
In the casino game keno, a machine randomly chooses \(20\) numbers between \(1\) and \(80\) (inclusive) without replacement. Players try to predict which numbers will be selected. Players don’t try to guess all \(20;\) generally, they’ll try to predict between \(1\) and \(10\) of the chosen numbers. The amount won depends on the number of guesses they made and the number of guesses that were correct. At one casino, a player can try to guess just one number. If that number is among the \(20\) selected, the player wins \($2;\) otherwise, the player loses \($1.\) What is the expected value?
- Answer
-
There are \(20\) winning numbers out of \(80\), so if we try to guess one of them, the probability of guessing correctly is \(\frac{20}{80}=\frac{1}{4}\). If the player guesses correctly, the player wins \(\$2\). The probability of losing is then \(\frac{3}{4}\). If the player guesses incorrectly, the player loses \(\$1\).
Outcomes and their corresponding Probabilities Outcomes (Events) Values \(n(O)\) Probability \(P(O)\) Winner \(\$2\) \(\frac{1}{4}\) Die \(-\$1\) \(\frac{3}{4}\) \begin{align*}
\text{Expected value} &=$2\times \frac{1}{4}+ (-$1)\times\frac{3}{4}\\
& = - $0.25\\
\end{align*}
By yourself or with a partner, devise your own lottery scheme. Assume you would have access to one or more machines that choose numbers randomly. What will a lottery draw look like? How many numbers are players choosing from? How many will be drawn? Will they be drawn with replacement or without replacement? What conditions must be met for a player to win the first or second (or more!) prize? Once you’ve decided on that, determine the payoff structure for winners and how much the game will cost. Try to make the game enticing enough that people will want to play it but with enough negative expected value that the lottery will make money. Aim for the expected value to be about \(−0.25\) times the cost of playing the game.
This page titled 3.5: Odd and Expected Value was last modified on Mon, 05 Oct 2026 18:04:01 GMT and is shared under a CC BY 4.0 license and was authored, remixed, and/or curated by Cheban Acharya via source content that was edited to the style and standards of the LibreTexts platform.
The book could not be loaded just now.